# Generative Retrieval with Few-shot Indexing

### v1

@@ 5 Conclusions
We have proposed a new, efficient, and effective GR paradigm, Few-Shot GR, featuring a novel few-shot indexing process that solely relies on prompting an LLM to record associations between queries and their docids, eliminating the need for any training steps. Given a query, Few-Shot GR conducts retrieval by promoting the LLM used for indexing and constraining it to generate a docid within the recorded docid created in indexing. We have designed few-shot indexing with one-to-many mapping to further enhance Few-Shot GR’s indexing. Experimental results demonstrate that GR achieves superior performance to state-of-the-art GR methods that require heavy training.
Limitations
We acknowledge the limitations of our work and outline avenues for future research. First, we only verify Few-Shot GR’s effectiveness on one dataset, NQ320k. It is valuable to investigate Few-Shot GR ’s performance on other ranking datasets, such as MS MARCO (Bajaj et al., 2016) and BEIR (Thakur et al., 2021). Second, existing work has shown that GR methods using training-based indexing perform worse as the corpus size increases (Pradeep et al., 2023). Since the dataset we used in our paper (NQ320K) has a document corpus with only 100k documents, we have yet to validate the effectiveness of Few-Shot GR on a document corpus with millions of documents. It is worthwhile to investigate whether Few-Shot GR’s effectiveness would generalise to a large-scale document corpus. Third, we claim that Few-Shot GR can potentially deal with a dynamic document corpus better than training-based indexing, because of Few-Shot GR’s non-training nature. This is because Few-Shot GR can easily add or remove docids in the docid bank created during few-shot indexing, and it does not suffer from the issue of forgetting. Several existing GR methods that use training-based indexing attempt to address this issue (Mehta et al., 2022; Kishore et al., 2023; Chen et al., 2023; Guo et al., 2024). It would be valuable to design experiments in the future to compare Few-Shot GR with these methods.
Acknowledgments
This research was partially supported by EU Horizon 2020 ITN/ETN on Domain Specific Systems for Information Extraction and Retrieval (H2020-EU.1.3.1., ID: 860721), and the China Scholarship Council (CSC) under grant number 202106220041.
Appendix A Appendix
Table 4: Case study of the 10 docids generated by Few-Shot GR for a document on NQ320K. Few-Shot GR uses Llama-3-8B-Instruct.
[…]
In accounting, minority interest (or non-controlling interest) is the portion of a subsidiary corporation’s stock that is not owned by the parent corporation. The magnitude of the minority interest in the subsidiary company is generally less than 50% of outstanding shares, or the corporation would generally cease to be a subsidiary of the parent. What is minority interest in accounting? minority-interest-accounting
What is non-controlling interest in accounting? non-controlling-interest-accounting
How is minority interest defined in accounting? minority-interest-definition
How is minority interest calculated in accounting? minority-interest-calculation
What is the significance of minority interest in accounting? minority-interest-significance
How does minority interest affect financial statements in accounting? minority-interest-financial-statements
How is minority interest treated in consolidated financial statements in accounting? minority-interest-consolidated-financial-statements
What is the impact of minority interest on the parent company’s earnings per share in accounting? minority-interest-impact-eps
How is minority interest reported in financial statements of a subsidiary company in accounting? minority-interest-reporting-subsidiary
What is the role of minority interest in a parent corporation’s financial statements? minority-interest-parent-corporation
Case study of docids generated by Few-Shot GR. Table A gives an example of 10 distinct docids generated by Few-Shot GR for a specific document in NQ320K. It shows that docids generated by Few-Shot GR are various.

### v2

@@ 5 Conclusions & Future Work
We have proposed a new, efficient, and effective GR paradigm, Few-Shot GR, featuring a few-shot indexing process that solely relies on prompting an LLM to record associations between queries and their docids, eliminating the need for any training steps. We have designed few-shot indexing with one-to-many mapping to further enhance Few-Shot GR’s indexing. Experimental results show that GR achieves superior performance to training-intensive state-of-the-art GR methods.
Suitability for dynamic corpora. Training-based indexing struggles with dynamic corpora, as training on new documents often causes forgetting of old ones [li2024matching]. Although several studies attempt to mitigate this issue [mehta2022dsi++, kishore2023incdsi, chen2023continual, guo2024corpusbrain++], it remains inherent to training-based methods. Few-Shot GR alleviates this challenge by enabling easy addition or removal of docids in the few-shot indexing docid bank, thus avoiding catastrophic forgetting. Future work can further explore this direction.
The datasets used in this paper, NQ320K and MS300K, contain corpora of 100K and 320K documents, respectively. So it is worthwhile to test whether Few-Shot GR’s effectiveness would generalise to a million-document corpus. Also, it is worth testing Few-Shot GR on other datasets (e.g., BEIR [thakur2021beir] and conversational search domains [mo2025conversational, meng2026conversational1, meng2025bridging, meng2023query]). Finally, exploring automatic retrieval quality prediction for generative retrieval methods is another promising direction [10.1145/3769733.3769743, meng2025query].